### v2

@@ 6 Our theory describes the behavior of deep networks on conjunction-wise additive tasks
So far, our analyses have been limited to simple kernel models, for their analytical tractability. We next test whether the insights developed from these simple models extend to a broader class of models: large-scale neural networks. We considered symbolic addition and context dependence on inputs created by concatenations of images from MNIST (Lecun et al., 1998) or CIFAR-10 (Krizhevsky et al., 2009). Each component corresponds to random samples from a particular category rather than the one-hot input considered so far and we study both in-distribution and compositional generalization. Notably, different image categories are not necessarily equally correlated with each other. To control for this, we randomly permuted the assignment of categories to components for each ($n=10$) experiment. We considered several relevant vision architectures: convolutional networks (ConvNets, LeCun et al., 1989) (trained on MNIST) and residual networks (ResNets, He et al., 2016) and Vision Transformers (ViTs, Dosovitskiy et al., 2020) (trained on CIFAR-10).
@@
Spatial distance of components impact salience in internal representations. First, we examined how changes in the input structure impact the networks’ representational geometry. We hypothesized that the ConvNets’ local weight structure should produce a more conjunctive representation for digits that are closer together. To test this, we determined $S(1;2)$ in an intermediate layer of the network, averaging over different instances of all digits. We found that $S(1;2)$ was indeed smaller for lower distances (Fig. 4a), indicating a more conjunctive population. This suggests that varying the distance between two digits provides a practical way of manipulating $S(1;2)$; below we use this insight to test predictions about how conjunctivity influences generalization behavior.
@@
Deep networks implement a conjunction-wise additive computation. Conjunction-wise additivity imposes a substantial computational restriction: on test set inputs, models can only add up values assigned to each conjunction of components seen during training. We therefore investigated whether the deep networks’ computations could be captured in these terms. Remarkably, we found that such a model was highly predictive of the model responses (see Section C.2). This indicates that large-scale neural networks tend to implement conjunction-wise additive computations — at least when trained on conjunction-wise additive tasks. Our theory suggests that they do so because the statistical structure of the dataset makes such a computation natural.
@@
Deep networks are impacted by a memorization leak on symbolic addition. Having found that our theory captures the networks’ general computational structure, we investigated whether it was also able to explain their specific performance. We first considered symbolic addition, training the ConvNets on seven different training sets with 20,000 samples and the ResNets and ViTs on three training sets with 40,000 samples. We varied both the size of $\mathcal{W}$ and whether generalization required interpolation or also extrapolation (Section D.1.3).
@@
Finally, our theory predicts that larger $\mathcal{W}$ should increase the compositional generalization slope. Our experiments confirmed this prediction (Fig. 4d). Further, whether the training set required interpolation or also extrapolation did not systematically affect the resulting slope, as predicted by our theory (though interestingly, it did affect in-distribution generalization). However, there was one exception to this: for $p=4$, the extrapolation dataset had a smaller slope than the interpolation data set. Our theory can therefore explain much but not all of the deep network behavior.
Deep networks are impacted by a shortcut bias on context dependence. Lastly, we trained the ConvNets on an MNIST version of context dependence using 30,000 training samples and the ResNets and ViTs on a CIFAR-10 version of the task using 40,000 training samples. Again, their behavior was aligned with the kernel theory’s predictions, having better-than-chance accuracy on CD-1 and CD-2, but having worse-than-chance accuracy on CD-3 (Fig. 4e). This confirms that the deep networks also suffered from a context-driven shortcut on CD-3.
7 Outlook: feature learning enables non-additive compositional generalization
In the previous section, we found that neural networks trained on symbolic addition or context dependence implement a conjunction-wise additive computation. However, conjunction-wise additivity is not the only mechanism by which deep neural networks generalize compositionally. To demonstrate this, we show in Appendix E that feature-learning networks can generalize on transitive equivalence (a non-additive task) due to their tendency to learn low-dimensional internal representations. This highlights that neural networks can implement non-additive compositional computations. Future work should explore more systematically how feature learning impacts compositional generalization.
@@ 8 Discussion
Humans often generalize to new situations by stitching together concepts and knowledge from prior experience in new ways. Despite the broad importance of this ability (both for cognitive science and machine learning), a general theory of when and how neural networks accomplish compositional generalization has remained elusive. Here we have taken a step towards formalizing this relationship by characterizing kernel models with compositionally structured representations, a framework that captures neural network training in the kernel regime and more broadly lets us understand the impact of dataset statistics on generalization. We found that they implement a specific compositional computation (“conjunction-wise additivity”). We then investigated how dataset statistics and representational geometry impact successful generalization, highlighting two failure modes arising from the inductive bias of gradient descent (memorization leak and shortcut bias). Finally, we validated our theory in deep neural networks trained on natural image data.
@@
Our results show how simple statistical models can implement (or at least approximate) abstract rules like context dependence. However, they also highlight that building in compositional structure by itself is often insufficient for compositional generalization. Indeed, contextual generalization is highly sensitive to the specific representational geometry and training data. This may explain why investigations into compositional generalization and the benefits of disentangled representation have yielded such inconsistent results. Overall, our insights highlight the utility of kernel models in systematically investigating model generalization (or at least generating an initial hypothesis as to the important factors). We here used these models to characterize deep neural networks, but they could be equally useful for better understanding human compositional generalization.
@@
While our work covers a broad range of different tasks, a number of limitations remain. Though out of scope here, it would be interesting to test this theory in extremely large-scale models, e.g. large language models. Future work should also consider inputs and outputs with broader ranges of formats (e.g. sequences with variable length) as well as representations that are not compositionally structured. Most importantly, our theory is limited to a particular learning mechanism (fixed features with linear readout); other learning mechanisms could either implement a conjunction-wise additive model that overcomes the failure modes highlighted here or implement novel compositional computations beyond conjunction-wise additivity (we gave an initial example of this in Section 7). By clarifying the general relationship between dataset statistics and compositional generalization, our theory provides an important foundation for such advances.
Acknowledgements
@@
We are grateful to the members of the Center for Theoretical Neuroscience for helpful comments and discussions. We thank Elom Amematsro, Ching Fang, and Matteo Alleman for detailed feedback. The work was supported by NSF 1707398 (Neuronex), Gatsby Charitable Foundation GAT3708, and the NSF AI Institute for Artificial and Natural Intelligence (ARNI).
@@ A.2 Why are we considering compositionally structured representations?
Characterizing compositional generalization in non-compositionally structured representations requires us to meaningfully distinguish between compositional and non-compositional aspects of their generalization. We consider this question out of scope for this manuscript, but believe it would be interesting future work.
@@ A.3 Nonlinear transformations and compositional structure
Further, the condition is met by most commonly used kernel functions including any radial basis function that depends on the Euclidean $\ell_{2}$-norm (e.g. the Gaussian kernel). Any nonlinear transform that meets this condition also conserves compositional structure, i.e. if $z$ is compositionally structured then so is $\phi(z)$. This allows us to derive computational restrictions on this broad range of models.
@@
Appendix B Representation salience: a metric for compositionally structured representations
@@ B.1 Definition
When emphasizing the total number of components, we denote $S(J;C)=S(J)$. Further, for compositionally structured representation, the similarity only depends on the total number of overlaps $|J|$. We therefore write $S(c;C)=S(J)$, where $c:=|J|$.
@@ B.4 Extended representational analysis
We computed the saliences by iteratively computing the representational similarities using the kernels derived in prior work (Cho & Saul, 2009; Tsuchida et al., 2018, 2019; Han et al., 2022). In Fig. 5b, we plot the different saliences for an input with four components. Now the salience of overlap 2 and 3 both first increase and then decrease and, just like for inputs with three components (Fig. 2), a Gaussian and rectified quadratic nonlinearity yields a particularly high salience for these intermediate conjunctions. Notably, they appear to be trading off the salience of these conjunctions differently: the rectified quadratic nonlinearity more strongly emphasizes overlaps of three whereas the Gaussian nonlinearity more strongly emphasizes overlaps of two. This highlights that for larger numbers of components, the dependence of generalization behavior on specific architectural choices will likely be even stronger.
@@
B.5 Proof of Proposition 5.1
See 5.1

### v3

@@ 6 Our theory can describe the behavior of deep networks on conjunction-wise additive tasks
So far, our analyses have been limited to simple kernel models, for their analytical tractability. We next tested whether the insights developed from these simple models extend to a broader class of models: large-scale neural networks. We considered symbolic addition and context dependence on inputs created by concatenations of images from MNIST (Lecun et al., 1998) or CIFAR-10 (Krizhevsky et al., 2009) (see Appendix D). Each component corresponds to random samples from a particular category rather than the one-hot input considered so far and we study both in-distribution and compositional generalization. Notably, different image categories are not necessarily equally correlated with each other. To control for this, we randomly permuted the assignment of categories to components for each ($n=10$) experiment. We considered several relevant vision architectures, which we all trained with backpropagation: convolutional networks (ConvNets, LeCun et al., 1989) (trained on MNIST) and residual networks (ResNets, He et al., 2016) and Vision Transformers (ViTs, Dosovitskiy et al., 2020) (trained on CIFAR-10).
@@
Spatial distance of components impacts salience in internal representations. First, we examined how changes in the input structure impact the networks’ representational geometry. We hypothesized that the ConvNets’ local weight structure should produce a more conjunctive representation for digits that are closer together. To test this, we determined $S(1;2)$ in an intermediate layer of the network, averaging over different instances of all digits. We found that $S(1;2)$ was indeed smaller for lower distances (Fig. 5a), indicating a more conjunctive population. This suggests that varying the distance between two digits provides a practical way of manipulating $S(1;2)$; below we use this insight to test predictions about how conjunctivity influences generalization behavior.
@@
Deep networks tend to implement conjunction-wise additive computations. Conjunction-wise additivity imposes a substantial computational restriction: on test set inputs, models can only add up values assigned to each conjunction of components seen during training. We therefore investigated whether the deep networks’ computations could be captured in these terms — a nontrivial prediction as these networks can in principle express arbitrary compositional computations. Remarkably, we found that a conjunction-wise additive model was highly predictive of the model responses (see Section D.3). This indicates that large-scale neural networks tend to implement conjunction-wise additive computations, at least when trained on conjunction-wise additive tasks. Our theory suggests that they do so because the statistical structure of the dataset makes such a computation natural.
@@
Deep networks are impacted by a memorization leak on symbolic addition. Having found that our theory captures the networks’ general computational structure, we investigated whether it was also able to explain their specific performance. We first considered symbolic addition, varying the size of the training set and whether generalization required interpolation or also extrapolation. We trained the ConvNets on 20,000 samples and the ResNets and ViTs on 40,000 samples (Section E.1.3).
@@
Finally, our theory predicts that larger training sets should increase the compositional generalization slope. Our experiments confirmed this prediction (Fig. 5d). Further, whether the training set required interpolation or also extrapolation did not systematically affect the resulting slope, as predicted by our theory (though interestingly, it did affect in-distribution generalization). Notably, there was one exception to the latter finding: for $p=4$, the extrapolation dataset had a smaller slope than the interpolation data set. Our theory can therefore explain much but not all of the deep network behavior.
Deep networks are impacted by a shortcut bias on context dependence. Lastly, we trained the ConvNets on an MNIST version of context dependence using 30,000 training samples and the ResNets and ViTs on a CIFAR-10 version of the task using 40,000 training samples. Again, their behavior was aligned with the kernel theory’s predictions, having better-than-chance accuracy on CD-1 and CD-2, but having worse-than-chance accuracy on CD-3 (Fig. 5e). This confirms that the deep networks also suffered from a context-driven shortcut on CD-3 (see also Fig. 16).
@@ 7 Discussion
Humans often generalize to new situations by stitching together concepts and knowledge from prior experience. Despite the broad importance of this ability (both for cognitive science and machine learning), a general theory of when and how neural networks accomplish compositional generalization has remained elusive. Here we have taken a step towards formalizing this relationship by characterizing kernel models with compositionally structured representations, a framework that captures neural network training or finetuning in the kernel regime and more broadly lets us understand the impact of representational geometry on generalization. We found that they implement a specific compositional computation (“conjunction-wise additivity”). We then investigated how dataset statistics and representational geometry impact successful generalization, highlighting two failure modes arising from the inductive bias of gradient descent (memorization leak and shortcut bias). Finally, we validated our theory in deep neural networks trained on natural image data.
@@
Our results show how simple statistical models can implement (or at least approximate) abstract rules like context dependence. However, they also highlight that building in compositional structure is often insufficient for compositional generalization. Indeed, contextual generalization is highly sensitive to the specific representational geometry and training data. This may explain why investigations into compositional generalization and the benefits of disentangled representations have yielded such inconsistent results. Overall, our insights highlight the utility of kernel models in systematically investigating model generalization. We here used these models to characterize deep neural networks, but they could be equally useful for better understanding human compositional generalization.
@@
While our work covers a broad range of different tasks, a number of limitations remain. We demonstrate that our theory captures many qualitative phenomena in deep neural networks, but do not provide any quantitative bounds. Further, though out of scope here, it would be interesting to test this theory in extremely large-scale models, e.g. large language models. Future work should also consider other input/output formats (e.g. sequences with variable length) and non-compositionally structured representations. Most importantly, our theory is limited to a particular learning mechanism (kernel models). Other learning mechanisms could overcome the limitations highlighted in Sections 4 and 5 (we give an initial example in Appendix F). By helping us design new benchmark tasks not solvable by kernel models, our theory provides an important foundation for such advances.
Acknowledgments
@@
We are grateful to the members of the Center for Theoretical Neuroscience for helpful comments and discussions. We thank Elom Amematsro, Ching Fang, and Matteo Alleman for detailed feedback, and Zeb Kurth-Nelson for comments on the manuscript. The work was supported by NSF 1707398 (Neuronex), Gatsby Charitable Foundation GAT3708, and the NSF AI Institute for Artificial and Natural Intelligence (ARNI).
@@ A.2 Why are we considering compositionally structured representations?
Characterizing compositional generalization in non-compositionally structured representations requires us to meaningfully distinguish between compositional and non-compositional aspects of their generalization. We present an example of such an analysis in Section A.5, proving that if the non-compositional representational components are random, the models still generalize in a conjunction-wise additive manner in expectation.
@@ A.3 Nonlinear transformations and compositional structure
Further, the condition is met by most commonly used kernel functions including any radial basis function that depends on the Euclidean $\ell_{2}$-norm (e.g. the Gaussian kernel). Any nonlinear transform that meets this condition also conserves compositional structure, i.e. if $x$ is compositionally structured then so is $\phi(x)$. This allows us to derive computational restrictions on this broad range of models.
@@
A.5 Non-compositionally structured representations
So far, we have considered representations that are compositionally structured. In practice, however, representations will almost never be exactly compositionally structured. Even when there is no specific similarity structure within components, there will be random noise in the representations. Here we characterize one such scenario, demonstrating that in expectation, these representations still yield conjunction-wise additive computations.
Proposition A.2.
Consider an input with $C$ components. We consider a representation that represents each component and conjunction of components $z_{J}$, $J\subseteq\{1,\dotsc,C\}$ by a random vector $x_{J}[z_{J}]\in\mathbb{R}^{d}$, $x_{J}\sim\mathcal{N}(0,\sigma_{k}^{2})$, $\sigma_{k}^{2}>0$, where $k=|J|$. The representation itself is given by the sum of all these vectors: $x=\sum_{J\subseteq\{1,\dotsc,C\}}x_{J}[z_{J}]$. All $x_{J}[z_{J}]$ are sampled independently from each other. This means that the similarity between different components as well as the entanglement of different components varies randomly. As a result, the kernel regression estimator $f(z)$ is not conjunction-wise additive. However, its expectation, $\mathbb{E}[f(z)]$, is conjunction-wise additive, indicating that all deviations from conjunction-wise additivity arise from random noise and cannot be used for systematic generalization.
Proof.
We consider the training representation $X=(x[z_{J}])_{z\in L^{\text{train}}}$, $X\in\mathbb{R}^{N_{\text{train}}\times d}$ and the test representation $\tilde{X}=(x[z_{J}])_{z\in L^{\text{test}}}$, $\tilde{X}\in\mathbb{R}^{N_{\text{test}}\times d}$. This gives rise to the training kernel $K=XX^{T}$ and the train-test kernel, $\tilde{K}=\tilde{X}X^{T}$. The dual coefficients are given by $a=K^{-1}y$. We now consider a test set input $\tilde{x}=\tilde{x}[\tilde{z}]$ representing the underlying components $\tilde{z}$. Denoting the similarity to the training set by $k(\tilde{x},X)=X\tilde{x}\in\mathbb{R}^{N_{\text{train}}}$, the test set behavior is given by
$f(\tilde{x})=a^{T}k(\tilde{x},X)=\sum_{J\in\text{Conj}(\tilde{z}|Z^{\text{train}})}a^{T}Xx_{J}[z_{J}]+\sum_{J\notin\text{Conj}(\tilde{z}|Z^{\text{train}})}a^{T}Xx_{J}[z_{J}]=:f_{1}(\tilde{x})+f_{2}(\tilde{x}),$ (11)
where we’ve simply split up the sum into the conjunction-wise additive part and the remainder. Clearly, this remainder, $f_{2}(\tilde{x}):=\sum_{J\notin\text{Conj}(\tilde{z}|Z^{\text{train}})}a^{T}Xx_{J}[z_{J}]$, will generally not be zero. However, we will now prove that $\mathbb{E}[f_{2}(\tilde{x})]=0$. As $f_{1}(\tilde{x})$ is conjunction-wise additive, this will prove the proposition. To do so, we note that we can assume that we have first sampled all Gaussian vectors relevant for the training set (we will denote this set of random variables by $\mathcal{X}^{\text{train}}$) and subsequently sample the set of Gaussian vectors only relevant for the test trial, i.e. $(\tilde{x}_{J})_{J\notin\text{Conj}(\tilde{z}|Z^{\text{train}})}$ (we will denote this set of random variables by $\mathcal{X}^{\text{test}}$). Then,
$\mathbb{E}[f_{2}(\tilde{x})]=\mathbb{E}_{\mathcal{X}^{\text{train}}}\left[\mathbb{E}_{\mathcal{X}^{\text{test}}}\left[f_{2}(\tilde{x})|\mathcal{X}^{\text{train}}\right]\right]=\sum_{J\notin\text{Conj}(\tilde{z}|Z^{\text{train}})}\mathbb{E}_{\mathcal{X}^{\text{train}}}\left[a^{T}X\mathbb{E}_{\mathcal{Z}^{\text{test}}}\left[x_{J}[\tilde{z}_{J}]|\mathcal{Z}^{\text{train}}\right]\right].$ (12)
This is zero, as $\mathbb{E}_{\mathcal{X}^{\text{test}}}\left[x_{J}[z_{J}]\right]=\mathbb{E}_{\mathcal{X}^{\text{test}}}\left[x_{J}[z_{J}]|\mathcal{X}^{\text{train}}\right]=0$, which follows from the fact that $x_{J}[z_{J}]$ for $J\notin\text{Conj}(\tilde{z}|Z^{\text{train}})$ is sampled independently from $\mathcal{X}^{\text{train}}$. ∎
Figure 6: Analysis of randomly sampled representations as considered in Proposition A.2. In all cases we sample 100-dimensional representations. For addition, we consider representations with weights that explore $S(1;2)\in[0.1,1]$ and for context dependence, we consider representations where $(\sigma_{1},\sigma_{2},\sigma_{3})\in\{(0.5,1,0.1),(1,0.5,0.1)\}$, where $\sigma_{k}$ indicates the standard deviation of the Gaussian vector representing conjunctions of $k$ components. a, We compare the additivity of individual models to the additivity of the model behavior averaged across 100 randomly sampled representations. While individual models may be severely non-additive, their average behavior is consistently highly additive, empirically confirming our proposition. b, As a result, when these models are trained on transitive equivalence, they also systematically exhibit chance performance, as they are still, on average, constrained to be conjunction-wise additive. c, We then estimated the slope of the model trained on symbolic addition for each random instance of a representation against corresponding representation’s $S(1;2)$. The individual slope estimates are depicted by the small translucent dots, whereas the average across all random instances is depicted by the larger points. Finally, our theoretical predictions in the exactly compositionally structured case are depicted by the lines. The average model behaviors are well described by our theory. d, We plot the distribution of generalization accuracies on context dependence across different model seeds, the two considered representations, and for CD-3 and CD-2. On CD-3, the representation with a higher salience for individual components (rep. 1, $(\sigma_{1},\sigma_{2},\sigma_{3})=(0.5,1,0.1)$ performs systematically below chance, whereas the representation with a higher salience for conjunctions of two components (rep. 2, $(\sigma_{1},\sigma_{2},\sigma_{3})=(1,0.5,0.1)$) generalizes above chance. In contrast, on CD-2, both representation exhibit better-than-chance-accuracy.
To test our theory empirically, we sample a range of Gaussian representations and train them on symbolic addition, transitive equivalence, and context dependence. For symbolic addition and transitive equivalence, we consider representations with expected $S(1;2)\in[0.1,1]$ using ten equally spaced values (i.e. $0.1,0.2,\dotsc$). For context dependence, we consider two types of representation: rep. 1, where $\sigma_{1}=1,\sigma_{2}=0.5,\sigma_{3}=0.1$, (i.e. single components are most saliently represented), and rep. 2, where $\sigma_{1}=0.5,\sigma_{2}=1,\sigma_{3}=0.1$ (i.e. conjunctions of two components are most saliently represented). We consider representations with $d=100$ and sample 100 instances of each representations. We then fit each representation to these three tasks.
We first estimate these models’ additivity on symbolic addition and context dependence. We find that while many model instance are highly additive, some can be highly non-additive (Fig. 6a). However, when averaging across the model behavior 100 randomly sampled representations, this average behavior is perfectly described by a conjunction-wise additive function. This empirically confirms our proposition.
Our proposition implies that these non-compositionally structured representations can still only systematically generalize on conjunction-wise additive tasks. To confirm this insight, we plotted the distribution of accuracies on the transitive equivalence problem across all random model instances. We found that while the models consistently learned the training set, they were indeed unable to generalize to the test set (Fig. 6b).
Finally, we investigated whether our analysis in Section 5 predicted the behavior of randomly sampled representations. Importantly, we have no theoretical guarantees for this scenario. Interestingly, however, the model behaviors were still well predicted by our theory. Specifically, we estimated the slopes that best described the compositional generalization behavior on symbolic addition for different training sets (as in Fig. 5). We found that across different model instances, these slopes were highly varied (Fig. 6c). However, the average slope across all random samples was well predicted by our analytical theory.
Similarly, our theoretical insights on context dependence provided meaningful insight into the generalization behavior of these randomly sampled models. Specifically, on CD-3, the model more saliently representing the single components failed to generalize, whereas the model more saliently representing conjunctions of two components generalized above chance. In contrast, on CD-2, both models generalized above chance. This indicates that our insights on how prevalent shortcut biases are for different training datasets extends to randomly sampled representations.
Appendix B Representational salience: a metric for compositionally structured representations
@@ B.1 Definition
When emphasizing the total number of components, we denote $S(J;C)=S(J)$. Further, for compositionally structured representation, the similarity only depends on the total number of overlaps $|J|$. We therefore write $S(k;C)=S(J)$, where $k:=|J|$.
To illustrate how salience would be computed in practice, we consider an example representation $X\in\mathbb{R}^{n\times d}$. This representation gives rise to a kernel $K=XX^{T}\in\mathbb{R}^{n\times n}$. We can also understand the trial-by-trial similarity in terms of the components it is overlapping in — denoting the set of all subsets of $\{1,\dotsc,C\}$ by $\mathcal{J}$, we denote this by $O\in\mathcal{J}^{n\times n}$. In particular, $O_{ii}=\{1,\dotsc,C\}$. For example, if data points $1$ and $2$ overlap in the third component, $O_{12}=\{3\}$. $\text{Sim}(J)$ is then defined as the entries $K_{ij}$ where $O_{ij}=J$. Note that in a compositionally structured representation, whenever $O_{ij}=J$, $K_{ij}$ takes on the same value, but more generally, we can define $\text{Sim}(J)$ by taking the average similarity. We can then define the unnormalized salience $\overline{S}$ by recursively computing it from $\text{Sim}(J)$ using Eq. 13. We then normalize the salience to get our final estimate.
@@ B.4 Extended representational analysis
We computed the saliences by iteratively computing the representational similarities using the kernels derived in prior work (Cho & Saul, 2009; Tsuchida et al., 2018; 2019; Han et al., 2022). In Fig. 7b, we plot the different saliences for an input with four components. Now the salience of overlap 2 and 3 both first increase and then decrease and, just like for inputs with three components (Fig. 3), a Gaussian and rectified quadratic nonlinearity yields a particularly high salience for these intermediate conjunctions. Notably, they appear to be trading off the salience of these conjunctions differently: the rectified quadratic nonlinearity more strongly emphasizes overlaps of three whereas the Gaussian nonlinearity more strongly emphasizes overlaps of two. This highlights that for larger numbers of components, the dependence of generalization behavior on specific architectural choices will likely be even stronger.
@@
B.5 Salience in very deep ReLU networks
Proposition B.2.